OpenStudy (anonymous):

rewrite the radical 240 in simplest form please

OpenStudy (anonymous):

sqrt 240 ? in simplest form?

OpenStudy (anonymous):

\[\sqrt{240}\] ?????

OpenStudy (anonymous):

= sqrt(16*15) = 4sqrt15

OpenStudy (anonymous):

i'm from uk so im not sure what radical means in this context - i believe its sqrt

OpenStudy (anonymous):

yes it's square and thanks for the correct answer.

OpenStudy (anonymous):

yw